A 2.0 kg block is sliding across a table with an initial speed of 0.5 m/s. The table imparts a 20.0 N force due to friction on the block. How much work in Joules will the table do on the block?
Answer #1

Work done by friction force will be equal to change in kinetic energy of block.

Wfriction = Kf - Ki​​​​​​

  = 0 - mu2/2

= 0 - 2x(0.52)/2

Wfriction = - 0.25 Joule